Prepare journal entries to record each of the following four separate issuances of stock.

  1. A corporation issued 8,000 shares of $20 par value common stock for $192,000 cash.
  2. A corporation issued 4,000 shares of no-par common stock to its promoters in exchange for their efforts, estimated to be worth $58,500. The stock has a $1 per share stated value.
  3. A corporation issued 4,000 shares of no-par common stock to its promoters in exchange for their efforts, estimated to be worth $58,500. The stock has no stated value.
  4. A corporation issued 2,000 shares of $50 par value preferred stock for $158,500 cash.

  Journal entry worksheet < A B C D Record the issue of 8,000 shares of $20 par value common stock for $192,000 cash. Note: Ent2) Record the issue of 4,000 shares of no-par common stock to its promoters in exchange for their efforts, estimated to be worth $58,500. The stock has a $1 per share stated value.

3) Record the issue of 4,000 shares of no-par common stock to its promoters in exchange for their efforts, estimated to be worth $58,500. The stock has no stated value.

4) Record the issue of 2,000 shares of $50 par value preferred stock for $158,500 cash.

Options For the Journal

  • Building
  • Cash
  • Common dividend payable
  • Common stock dividend distributable
  • Common stock, $1 par value
  • Common stock, $1 stated value
  • Common stock, $20 par value
  • Common stock, no-par value
  • Contributed capital, treasury stock
  • Income summary
  • Inventory
  • Land
  • Machinery
  • Note payable
  • Organization expenses
  • Paid-in capital in excess of par value, Common stock
  • Paid-in capital in excess of par value, preferred stock
  • Paid-in capital in excess of stated value, common stock
  • Preferred stock, $0.50 par value
  • Preferred stock, $1 stated value
  • Preferred stock, $1 stated value
  • Preferred stock, $20 par value
  • Preferred stock, $5 par value
  • Preferred stock, $50 par value
  • Preferred stock, no-par value
Answer #1

Answer-

Journal Entries
DATE GENERAL JOURNAL DEBIT CREDIT
1 Cash 192,000
Common Stock (8,000*20) 160,000
Paid in Excess of Par- Common Stock = 192,000-160,000 32,000
(To record common shares isused for cash)
2 Promotion Expenses 58,500
Common Stock (4,000*1) 4,000
Paid in Excess of Par- Common Stock = 58,500-4,000 54,500
(To record common shares isused to promoters)
3 Promotion Expenses 58,500
Common Stock 58,500
(To record 4,000 common shares isused to promoters)
4 Cash 158,500
Preferred Stock (2,000*50) 100,000
Paid in Excess of Par- preferred Stock = 158,500-100,000 58,500
(To record preferred shares isused for cash)

Kindly give Thumbs up if u like my answer....Thanks!!!!